# Linear colorings of graphs

### v1

@@ 4.3 Linearly coloring caterpillars
Let $P$ be the central path of $T$, and $k=\lceil\log(|V(P)|+1)\rceil$. We know from Section 2.2 that $\operatorname{\chi_{\mathrm{cen}}}(P)=\operatorname{\chi_{\mathrm{lin}}}(P)=k$. We define a coloring for $T$ as follows: the path $P$ is colored according to a centered coloring achieving $\operatorname{\chi_{\mathrm{cen}}}(P)=k$ and all the leaves of $T$ receive a $(k+1)$st color. This is a centered coloring of $T$ as each connected subgraph $T^{\prime}$ of $T$ that has more than one vertex is composed of a nonempty subpath of $P$, which has a center, and a (potentially empty) subset of leaves of $T$ that do not share any color with the subpath. Thus, the center of the subpath is also a center of $T^{\prime}$. Therefore, $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant k+1$. Moreover, $T$ contains $P$ with $\operatorname{\chi_{\mathrm{lin}}}(P)=k$ as a subgraph so by Section 2.2, $\operatorname{\chi_{\mathrm{lin}}}(T)\geqslant k$. Hence $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant\operatorname{\chi_{\mathrm{lin}}}(T)+1$. ∎
@@ 5 Obstructions to bounded linear chromatic number
[[NOdM12]] For every $k$ the class of graphs of centered chromatic number at most $k$ is well-quasi-ordered by the induced subgraph relation.
@@ 5.1 Small values
A graph $G$ satisfies $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 1$ if and only if $G$ is edgeless, and $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$ if and only if $G$ is star forest. In both cases $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@
Towards a contradiction, denote such a cycle by $C$, and let $v_{1},v_{2},v_{3},v_{4}$ be a cyclic order of the vertices of $C$ such that $v_{2}$ and $v_{4}$ have degree $2$ in $G$. We first show that there exists some $i\in\{1,3\}$ such that $v_{i}$ is not an endpoint of a path of length $2$ in the graph $G-v_{4-i}$. For $i\in\{1,3\}$, let $X_{i}$ be the set of vertices of $G$ adjacent to $v_{i}$ but not to $v_{4-i}$. Then, $X_{1}$ and $X_{3}$ are disjoint (by definition) and there is no path from $X_{1}$ to $X_{3}$ in the graph $G-v_{1}-v_{3}$, since otherwise $G$ would contain a cycle of length more than $4$. Suppose for a contradiction that there is a path of length $2$, say $v_{1}x_{1}y_{1}$, in the graph $G-v_{3}$ and a path of length $2$, say $v_{3}x_{3}y_{3}$, in the graph $G-v_{1}$. Since vertices $v_{2}$ and $v_{4}$ have degree $2$ in $G$, they are both distinct from $x_{1}$ and $x_{3}$, as well as from $y_{1}$ and $y_{3}$. Note that no vertex adjacent to both $v_{1}$ and $v_{3}$ can be adjacent to a vertex outside of $C$, otherwise we would obtain an $F_{9}$ subgraph. In particular, we have $x_{i}\in X_{i}$ for $i\in\{1,3\}$. Then, the graph $G$ contains three edge-disjoint paths of lengths $1$, $2$, and $4$ starting at $v_{1}$, namely $v_{1}v_{2}$, $v_{1}x_{1}y_{1}$, and $v_{1}v_{4}v_{3}x_{3}y_{3}$, respectively. These paths form a subgraph of $G$ isomorphic to $F_{5}\in\mathcal{F}$, a contradiction. This shows that there exists some $i\in\{1,3\}$ such that $v_{i}$ is not an endpoint of a path of length $2$ in the graph $G-v_{4-i}$.
@@
We may assume without loss of generality that $v_{3}$ is the center of a star component of the graph $G-v_{1}$. We claim that the graph $G-v_{1}$ is a star forest. Suppose for a contradiction that this is not the case. Suppose first that $G-v_{1}$ contains a cycle, say $C^{\prime}$. Then $C^{\prime}$ does not belong to the component of $G-v_{1}$ containing $v_{3}$ and we conclude that the graph $G-v_{3}$ contains a path from $v_{1}$ to $V(C^{\prime})$; let $P$ be a shortest such path. Note that $P$ is of length one, since otherwise the path obtained by the concatenation of the paths $v_{4}v_{3}v_{2}v_{1}$, $P$, and a path in $C^{\prime}$ of length $2$ would form a path of length at least $7$ in $G$, a contradiction with the fact that $G$ does not contain $P_{8}$ as a subgraph. A similar argument shows that $C^{\prime}$ is the $3$-cycle. But now, $G$ contains $F_{6}$ as a subgraph, a contradiction. This shows that $G-v_{1}$ is acyclic. Since $G-v_{1}$ is not a star forest, it contains a path of length $3$ as a subgraph; let $Q=w_{1}w_{2}w_{3}w_{4}$ be such a path. As above, $Q$ does not belong to the component of $G-v_{1}$ containing $v_{3}$, hence, the graph $G-v_{3}$ contains a path from $v_{1}$ to $V(Q)$; let $R$ be a shortest such path. Similarly as above, we first observe that $R$ has length one; furthermore, since $G$ does not contain $P_{8}$ as a subgraph, $R$ must attach to an internal vertex of $Q$, say $w_{2}$. But now, $G$ contains three edge-disjoint paths of lengths $1$, $2$, and $4$ starting at $w_{2}$, namely $w_{2}w_{1}$, $w_{2}w_{3}w_{4}$, and $w_{2}v_{1}v_{2}v_{3}v_{4}$, respectively, forming a subgraph of $G$ isomorphic to $F_{5}\in\mathcal{F}$, a contradiction. This shows that $G-v_{1}$ is a star forest.
@@
Suppose for a contradiction that $G$ contains a $2$-connected subgraph $G^{\prime}$ on at least $5$ vertices. Then $G^{\prime}$ has to contain a cycle on $4$ vertices, say $C$, with vertices $v_{1},v_{2},v_{3},v_{4}$ in cyclic order. Let $v_{5}$ be a vertex of $V(G^{\prime})\setminus V(C)$. Then, by $2$-connectivity, there are two paths in $G^{\prime}$ connecting $v_{5}$ to $C$ that are vertex-disjoint except in $v_{5}$. Furthermore, each of these two paths has length one, since otherwise $G^{\prime}$ would contain a cycle of length greater than $4$. For the same reason, these two paths connect to $C$ at opposite vertices, with loss of generality at $v_{1}$ and $v_{3}$. Notice that $G^{\prime}$ does not contain an edge with endpoints in $\{v_{2},v_{4},v_{5}\}$, as otherwise $G^{\prime}$ contains a $5$-cycle with vertex set $V(C)\cup\{v_{5}\}$. Furthermore, none of the vertices in $\{v_{2},v_{4},v_{5}\}$ is adjacent in $G$ to a vertex outside $V(C)$, since otherwise $G$ would contain $F_{9}$ as a subgraph. This implies that $v_{2}$ and $v_{4}$ are two non-adjacent vertices of $C$ that have degree $2$ in $G$, contradicting Section 5.1. ∎
@@
Suppose for a contradiction that $G$ is a forest. Since $G$ is connected, $G$ is a tree. The minimality of $G$ also implies that $G$ does not have any pairs of twin leaves: indeed, if $u$ and $v$ are two leaves of $G$ with the same neighbor, then any linear $3$-coloring of $G-v$ can be extended to a linear $3$-coloring of $G$ by assigning to $v$ the color of $u$, contradicting the fact that $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant 4$. Let $P=v_{1}\dots v_{k}$ be a longest path in $G$. Then $k\leqslant 7$ since $G$ contains no $P_{8}$ and $k\geqslant 4$ since otherwise $G$ is a star, implying by Section 5.1 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$, a contradiction. By the maximality of $P$, $d(v_{1})=d(v_{k})=1$. By the absence of twin leaves, $d(v_{2})=d(v_{k-1})=2$. We consider three cases depending on the value of $k$.
@@
    Case $k\in\{4,5\}$. In this case, the maximality of $P$ implies that every vertex of $G$ is at distance at most $2$ from $v_{3}$. Therefore, deleting $v_{3}$ from $G$ results in star forest, implying by Sections 5.1 and 2.2 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 3$, a contradiction.
@@
    Case $k=6$. Only $v_{3}$ and $v_{4}$ may have neighbors outside of $P$. However, by the maximality of $P$, none of $v_{3}$ and $v_{4}$ can be an endpoint of a path of length more than $2$ that is edge-disjoint from $P$. Moreover, $v_{3}$ and $v_{4}$ cannot both be endpoints of paths of length $2$ that are edge-disjoint from $P$ without creating a $F_{4}$ subgraph. We can thus assume without loss of generality that every path with endpoint $v_{4}$ that is edge-disjoint from $P$ has length at most $1$. Moreover, since $G$ has no twin leaves, there can be only one such leaf $v$. But now, $G$ admits a linear coloring with three colors as follows. Vertex $v_{5}$ gets color $1$, vertices $v_{4}$ and $v_{6}$ get color $2$, vertex $v_{3}$ gets color $3$, and so does $v$ (if it exists). Then, for the remaining vertices, we color all vertices at distance $i$ from $v_{3}$ with color $i$ for $i\in\{1,2\}$. It is easy to check that this is indeed a linear $3$-coloring of $G$, a contradiction.
@@
    Case $k=7$. As $G$ does not have $F_{5}$ as subgraph, $d(v_{3})=d(v_{5})=2$. So $v_{4}$ is the only vertex of $P$ that can have degree more than $2$. Suppose that $v_{4}$ has a neighbor $v$ that does not belong to $P$. Then, by the maximality of $P$, the vertex $v$ cannot be an endpoint of a path of length more than $2$ in the graph $G-v_{4}$. Furthermore, if $v$ is an endpoint of such a path of length $2$, then $v$ has degree $2$ in $G$, since otherwise $G$ would contain $F_{5}$ as a subgraph. Hence, each component of $G-v_{4}$ is a star, implying by Sections 5.1 and 2.2 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 3$, a contradiction. ∎
@@
Suppose for a contradiction that $G$ contains at least two copies of $K_{3}$. Then, at least one of those copies is disjoint from $P$. Since $v_{1}$ is adjacent to a vertex from $P$, this gives rise to either $F_{6}$ or $F_{7}$, a contradiction. ∎
As in the proof of Section 5.1, for every vertex $v\in C$, the graph $G-v$ contains a $P_{4}$ adjacent to $v$. We say that $v\in V(C)$ is good if the unique edge in $C-v$ is the middle edge of all the copies of $P_{4}$ in $G-v$.
@@
The above claim implies that there is at most one good vertex of $C$. To conclude the proof it is enough to observe that, in the case of a single good vertex we obtain an $F_{2}$, while in the case of no good vertex we obtain an $F_{8}$, a contradiction.
@@ 6 Algorithmic aspects
We now discuss the complexity of Linear Chromatic Number, the decision problem that takes as input a graph $G$ and an integer $k$, and the task is to determine if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$. It is not clear if the problem is in NP, since, as shown by Kun et al. [KOPS21], the problem of determining whether a coloring is linear is co-NP-complete. Nevertheless, the close relationship between the centered and linear chromatic numbers together with known hardness results for the centered chromatic number lead to the following result. A graph $G$ is cobipartite if its vertex set is a union of two cliques.
@@
Every connected cobipartite graph has a Hamiltonian path and every connected induced subgraph of a cobipartite graph is cobipartite. Thus, by Section 4.1, any linear coloring of a connected cobipartite graph $G$ is also a centered coloring, hence, $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\chi_{\mathrm{lin}}}(G)$. This implies that Linear Chromatic Number, when restricted to cobipartite graphs, is in NP: if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$, then this can be certified by giving a treedepth decomposition with depth at most $k$, which is an equivalent definition of centered chromatic number (see [NOdM12]). As shown by Bodlaender et al. [BDJ+98], given a cobipartite graph $G$ and an integer $k$, it is NP-complete to determine if $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant k$. Since $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant k$ if and only if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$, the claimed NP-completeness follows. ∎

### v2

@@ 4.3 Linearly coloring caterpillars
Let $P$ be the central path of $T$, and $k=\lceil\log(|V(P)|+1)\rceil$. We know from Section 2.2 that $\operatorname{\chi_{\mathrm{cen}}}(P)=\operatorname{\chi_{\mathrm{lin}}}(P)=k$. We define a coloring for $T$ as follows: the path $P$ is colored according to a centered coloring achieving $\operatorname{\chi_{\mathrm{cen}}}(P)=k$ and all the leaves of $T$ receive a (k+1)(k+1)st color. This is a centered coloring of $T$ as each connected subgraph $T^{\prime}$ of $T$ that has more than one vertex is composed of a nonempty subpath of $P$, which has a center, and a (potentially empty) subset of leaves of $T$ that do not share any color with the subpath. Thus, the center of the subpath is also a center of $T^{\prime}$. Therefore, $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant k+1$. Moreover, $T$ contains $P$ with $\operatorname{\chi_{\mathrm{lin}}}(P)=k$ as a subgraph so by Section 2.2, $\operatorname{\chi_{\mathrm{lin}}}(T)\geqslant k$. Hence, $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant\operatorname{\chi_{\mathrm{lin}}}(T)+1$. ∎
@@ 5 Obstructions to bounded linear chromatic number
[[NO12]] For every $k$ the class of graphs of centered chromatic number at most $k$ is well-quasi-ordered by the induced subgraph relation.
@@ 5.1 Small values
A graph $G$ satisfies $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 1$ if and only if $G$ is edgeless, and $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$ if and only if $G$ is a star forest. In both cases $\operatorname{\chi_{\mathrm{lin}}}(G)=\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@
Towards a contradiction, denote such a cycle by $C$, and let $v_{1},v_{2},v_{3},v_{4}$ be a cyclic order of the vertices of $C$ such that $v_{2}$ and $v_{4}$ have degree $2$ in $G$. We first show that there exists some $i\in\{1,3\}$ such that $v_{i}$ is not an endpoint of a path of length $2$ in the graph $G-v_{4-i}$. For $i\in\{1,3\}$, let $X_{i}$ be the set of vertices of $G$ adjacent to $v_{i}$ but not to $v_{4-i}$. Then, $X_{1}$ and $X_{3}$ are disjoint (by definition) and there is no path from $X_{1}$ to $X_{3}$ in the graph $G-v_{1}-v_{3}$, since otherwise $G$ would contain a cycle of length more than $4$. Suppose for a contradiction that there is a path of length $2$, say $v_{1}x_{1}y_{1}$, in the graph $G-v_{3}$ and a path of length $2$, say $v_{3}x_{3}y_{3}$, in the graph $G-v_{1}$. Since vertices $v_{2}$ and $v_{4}$ have degree $2$ in $G$, they are both distinct from $x_{1}$ and $x_{3}$, as well as from $y_{1}$ and $y_{3}$. Also, $y_{1}$ and $y_{3}$ are distinct, since there is no path from $X_{1}$ to $X_{3}$ in $G-v_{1}-v_{3}$. Note that no vertex adjacent to both $v_{1}$ and $v_{3}$ can be adjacent to a vertex outside of $C$, otherwise we would obtain an $F_{9}$ subgraph. In particular, we have $x_{i}\in X_{i}$ for $i\in\{1,3\}$. Then, the graph $G$ contains three edge-disjoint paths of lengths $1$, 22, and $4$ starting at $v_{1}$, namely $v_{1}v_{2}$, $v_{1}x_{1}y_{1}$, and $v_{1}v_{4}v_{3}x_{3}y_{3}$, respectively. These paths form a subgraph of $G$ isomorphic to $F_{5}\in\mathcal{F}$, a contradiction. This shows that there exists some $i\in\{1,3\}$ such that $v_{i}$ is not an endpoint of a path of length $2$ in the graph $G-v_{4-i}$.
@@
We may assume without loss of generality that $v_{3}$ is the center of a star component of the graph $G-v_{1}$. We claim that the graph $G-v_{1}$ is a star forest. Suppose for a contradiction that this is not the case. Suppose first that $G-v_{1}$ contains a cycle, say $C^{\prime}$. Then $C^{\prime}$ does not belong to the component of $G-v_{1}$ containing $v_{3}$ and we conclude that the graph $G-v_{3}$ contains a path from $v_{1}$ to $V(C^{\prime})$; let $P$ be a shortest such path. Note that $P$ is of length one, since otherwise the path obtained by the concatenation of the paths $v_{4}v_{3}v_{2}v_{1}$, $P$, and a path in $C^{\prime}$ of length $2$ would form a path of length at least $7$ in $G$, a contradiction with the fact that $G$ does not contain $P_{8}$ as a subgraph. A similar argument shows that $C^{\prime}$ is the $3$-cycle. But now, $G$ contains $F_{6}$ as a subgraph, a contradiction. This shows that $G-v_{1}$ is acyclic. Since $G-v_{1}$ is not a star forest by assumption, it contains a path of length $3$ as a subgraph; let $Q=w_{1}w_{2}w_{3}w_{4}$ be such a path. As above, $Q$ does not belong to the component of $G-v_{1}$ containing $v_{3}$, hence, the graph $G-v_{3}$ contains a path from $v_{1}$ to $V(Q)$; let $R$ be a shortest such path. Similarly as above, we first observe that $R$ has length one; furthermore, since $G$ does not contain $P_{8}$ as a subgraph, $R$ must attach to an internal vertex of $Q$, say $w_{2}$. But now, $G$ contains three edge-disjoint paths of lengths $1$, 22, and $4$ starting at $w_{2}$, namely $w_{2}w_{1}$, $w_{2}w_{3}w_{4}$, and $w_{2}v_{1}v_{2}v_{3}v_{4}$, respectively, forming a subgraph of $G$ isomorphic to $F_{5}\in\mathcal{F}$, a contradiction. This shows that $G-v_{1}$ is a star forest.
@@
Suppose for a contradiction that $G$ contains a $2$-connected subgraph $G^{\prime}$ on at least $5$ vertices. Then $G^{\prime}$ has to contain a cycle on $4$ vertices, say $C$, with vertices $v_{1},v_{2},v_{3},v_{4}$ in cyclic order. Let $v_{5}$ be a vertex of $V(G^{\prime})\setminus V(C)$. Then, by $2$-connectivity, there are two paths in $G^{\prime}$ connecting $v_{5}$ to $C$ that are vertex-disjoint except in $v_{5}$. Furthermore, each of these two paths has length one, since otherwise $G^{\prime}$ would contain a cycle of length greater than $4$. For the same reason, these two paths connect to $C$ at opposite vertices, with loss of generality at $v_{1}$ and $v_{3}$. Notice that $G^{\prime}$ does not contain an edge with endpoints in $\{v_{2},v_{4},v_{5}\}$, as otherwise $G^{\prime}$ would contain a $5$-cycle with vertex set $V(C)\cup\{v_{5}\}$. Furthermore, none of the vertices in $\{v_{2},v_{4},v_{5}\}$ is adjacent in $G$ to a vertex outside $V(C)$, since otherwise $G$ would contain $F_{9}$ as a subgraph. This implies that $v_{2}$ and $v_{4}$ are two non-adjacent vertices of $C$ that have degree $2$ in $G$, contradicting Section 5.1. ∎
@@
Suppose for a contradiction that $G$ is a forest. Since $G$ is connected, $G$ is a tree. The minimality of $G$ also implies that $G$ does not have any pairs of twin leaves: indeed, if $u$ and $v$ are two leaves of $G$ with the same neighbor, then any linear $3$-coloring of GvG-v can be extended to a linear $3$-coloring of $G$ by assigning to $v$ the color of $u$, contradicting the fact that $\operatorname{\chi_{\mathrm{lin}}}(G)\geqslant 4$. Let $P=v_{1}\dots v_{k}$ be a longest path in $G$. Then $k\leqslant 7$ since $G$ contains no $P_{8}$ and $k\geqslant 4$ since otherwise $G$ is a star, implying by Section 5.1 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 2$, a contradiction. By the maximality of $P$, $d(v_{1})=d(v_{k})=1$. By the absence of twin leaves, $d(v_{2})=d(v_{k-1})=2$. We consider three cases depending on the value of $k$.
@@
    Case $k\in\{4,5\}$. In this case, the maximality of $P$ implies that every vertex of $G$ is at distance at most $2$ from $v_{3}$. Therefore, deleting $v_{3}$ from $G$ results in a star forest, implying by Sections 5.1 and 2.2 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 3$, a contradiction.
@@
    Case $k=6$. Only $v_{3}$ and $v_{4}$ may have neighbors outside of $P$. However, by the maximality of $P$, neither of $v_{3}$ and $v_{4}$ can be an endpoint of a path of length more than $2$ that is edge-disjoint from $P$. Moreover, $v_{3}$ and $v_{4}$ cannot both be endpoints of paths of length $2$ that are edge-disjoint from $P$ without creating an $F_{4}$ subgraph. We can thus assume without loss of generality that every path with endpoint $v_{4}$ that is edge-disjoint from $P$ has length at most $1$. Moreover, since $G$ has no twin leaves, there can be only one such leaf $v$. But now, $G$ admits a linear coloring with three colors as follows. Vertex $v_{5}$ gets color $1$, vertices $v_{4}$ and $v_{6}$ get color $2$, vertex $v_{3}$ gets color $3$, and so does $v$ (if it exists). Then, for the remaining vertices, we color all vertices at distance $i$ from $v_{3}$ with color $i$ for $i\in\{1,2\}$. It is easy to check that this is indeed a linear $3$-coloring of $G$, a contradiction.
@@
    Case $k=7$. As $G$ does not have $F_{5}$ as a subgraph, $d(v_{3})=d(v_{5})=2$. So $v_{4}$ is the only vertex of $P$ that can have degree more than $2$. Suppose that $v_{4}$ has a neighbor $v$ that does not belong to $P$. Then, by the maximality of $P$, the vertex $v$ cannot be an endpoint of a path of length more than $2$ in the graph $G-v_{4}$. Furthermore, if $v$ is an endpoint of such a path of length $2$, then $v$ has degree $2$ in $G$, since otherwise $G$ would contain $F_{5}$ as a subgraph. Hence, each component of $G-v_{4}$ is a star, implying by Sections 5.1 and 2.2 that $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant 3$, a contradiction. ∎
@@
Suppose for a contradiction that $G$ contains a second copy $C^{\prime}$ of $K_{3}$ (i.e., $C^{\prime}\neq C$, but they may share vertices or edges). As noted above, $C^{\prime}$ contains $v_{1}$. Also, $C$ and $C^{\prime}$ cannot share other vertices because of Section 5.1 and the fact that $K_{4}$ belongs to $\mathcal{F}$. Observe that $P$ does not intersect both $C$ and $C^{\prime}$, as that would result in a longer cycle. Thus, either $C$ or $C^{\prime}$ is disjoint from $P$. Since $v_{1}$ is adjacent to a vertex from $P$, this gives rise to either $F_{6}$ or $F_{7}$, a contradiction. ∎
As in the proof of Section 5.1, for every vertex $v\in C$, the graph GvG-v contains a $P_{4}$ adjacent to $v$. We say that $v\in V(C)$ is good if the unique edge in CvC-v is the middle edge of all the copies of $P_{4}$ in GvG-v.
@@
The above claim implies that there is at most one good vertex of $C$. To conclude the proof it is enough to observe that, in the case of a single good vertex, we obtain an $F_{2}$, while in the case of no good vertex we obtain an $F_{8}$, a contradiction.
@@ 6 Algorithmic aspects
We now discuss the complexity of Linear Chromatic Number, the decision problem that takes as input a graph $G$ and an integer $k$, and the task is to determine if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$. It is not clear if the problem is in NP, since, as shown by Kun et al. [KOP+21], the problem of determining whether a coloring is linear is co-NP-complete. Nevertheless, the close relationship between the centered and linear chromatic numbers, together with known hardness results for the centered chromatic number, lead to the following result. A graph $G$ is cobipartite if its vertex set is a union of two cliques.
@@
Every connected cobipartite graph has a Hamiltonian path and every connected induced subgraph of a cobipartite graph is cobipartite. Thus, by Section 4.1, any linear coloring of a connected cobipartite graph $G$ is also a centered coloring, hence, $\operatorname{\chi_{\mathrm{cen}}}(G)=\operatorname{\chi_{\mathrm{lin}}}(G)$. This implies that Linear Chromatic Number, when restricted to cobipartite graphs, is in NP: if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$, then this can be certified by giving a treedepth decomposition with depth at most $k$, which is an equivalent definition of centered chromatic number (see [NO12]). As shown by Bodlaender et al. [BDJ+98b], given a cobipartite graph $G$ and an integer $k$, it is NP-complete to determine if $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant k$. Since $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant k$ if and only if $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant k$, the claimed NP-completeness follows. ∎